Astrid1:

@sNoWfLaKe0541 Mahth help.... Be ready for me being dumb. <3 *uploading ss*

Astrid1:

Astrid1:

@snowflake0531

Astrid1:

It's the The Pythagorean Theorem. e.e Idk how to do this.

snowflake0531:

In a right triangle, where one of the angles is 90 degrees a and b ar ehte legs of the triangle, while c is the hypotenuse \[a^2 + b^2=c^2\]

snowflake0531:

In triangle a, we have the lengths of a and b, can you put it into teh equation?

snowflake0531:

\[a^2 + b^2\] a is 2, b is 4 what is it

Astrid1:

\[2^2+4^2= 20?\]

snowflake0531:

Yep so \[20=c^2\] c equals what

LouisAKALuis:

ohhhhh

Astrid1:

\[ 400?\]

snowflake0531:

20^2 = c c would equal 400 But i'm saying 20 equals c squared

snowflake0531:

\[20=c^2\] \[\sqrt20 = \sqrt c^2\]

snowflake0531:

So, \[c=\sqrt20\]

Astrid1:

I just looked at my math book.. It shows somethng quite like this.. So my answer would be 10? or maybe 5? Or am Ijust plain stupid🤦

snowflake0531:

fhdjglKUZ<>%^E$WTr56ey7tdr6fewsdgrxzc4a plain stupid

Astrid1:

Yeh... So uhh...

snowflake0531:

do you know how sqrt stuff works?..

snowflake0531:

It either stays as it is, or you find a perfect square in it, in this case there's a perfect square in it

snowflake0531:

\[\sqrt20 = \sqrt(4\times5) = 2\sqrt5\]

snowflake0531:

do you agree, that 4 times 5 equals 20

Astrid1:

Yea Yea Yeaaa.

snowflake0531:

so we have \[\sqrt (4 \times 5) = \sqrt4 \times \sqrt5\] agreed?

Astrid1:

Yeah..

snowflake0531:

so what is sqrt4

Astrid1:

2?

snowflake0531:

yep so what is 2 times sqrt5 is 2sqrt5

Astrid1:

\[2 \times \sqrt{5}\] uhmmmm...... uhmmmmmm.......... After a long pause she figured she may as well thinkg.. 4.472135955?

snowflake0531:

e.e i think you can make it stay at 2sqrt5 lol

Astrid1:

Ok...